# COEF-VQ: Cost-Efficient Video Quality Understanding through a Cascaded Multimodal LLM Framework

### v1

@@ Abstract.
Recently, with the emergence of recent Multimodal Large Language Model (MLLM) technology, it has become possible to exploit its video understanding capability on different classification tasks. In practice, we face the difficulty of huge requirements for GPU resource if we need to deploy MLLMs online. In this paper, we propose COEF-VQ, a novel cascaded MLLM framework for better video quality understanding on TikTok. To this end, we first propose a MLLM fusing all visual, textual and audio signals, and then develop a cascade framework with a lightweight model as pre-filtering stage and MLLM as fine-consideration stage, significantly reducing the need for GPU resource, while retaining the performance demonstrated solely by MLLM. To demonstrate the effectiveness of COEF-VQ, we deployed this new framework onto the video management platform (VMP) at TikTok, and performed a series of detailed experiments on two in-house tasks related to video quality understanding. We show that COEF-VQ leads to substantial performance gains with limit resource consumption in these two tasks.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systemsccs: Computing methodologies Artificial intelligenceccs: Information systems Language models
@@ 1. Introduction
Figure 1. Illustration of our proposed unified Multimodal LLM architecture: video frames and text tokens are processed and early-fused through a VL foundational model, and audio signal is processed through an Audio encoder and then late-fused with the hidden representations from VL model. The MLLM will be SFT-ed on domain dataset before deployment.
@@
Therefore, in this paper we propose COEF-VQ, which is composed of two key novelties: 1)1)1 ) we first design a unified Multimodal LLM framework integrating video frames, texts, and audio information, which goes beyond the existing multi-tower baseline deployed on our platform (Conneau, 2019; He et al., 2016; Liu et al., 2021; Radford et al., 2023) and typical industry practices (Zhai et al., 2023). This enhanced model architecture allows for more sophisticated feature fusion and deeper cross-modal interactions, boosting the video classification performance. By doing so, it alleviates the limitations of traditional multi-tower structures, making the model more sensitive and accurate in discerning domain-specific tasks.
@@
However, despite the powerful capabilities of MLLM, the large number of parameters in MLLM, which enables them to perform robust multimodal reasoning, requires substantial GPU memory and processing speed. This demand, when scaled for online deployment on platforms like TikTok, would lead to considerable GPU resource consumption, affecting both deployment costs and latency in real-time, online video service.
2)2)2 ) we develop a cascade serving structure for online deployment, designed to optimize resource usage with minimal performance sacrifice. In the first stage, predictions are made using a lightweight model (usually the baseline) at the time of video posting. Based on the lightweight model’s prediction or the number of video views (VV), videos could be considerably filtered, reducing the volume entering the next stage. In the second stage, MLLM predicts the videos passing through the first stage with much higher accuracy, significantly reducing false positives. The proposed cascade structure offers an efficient, scalable solution to the deployment challenges of MLLMs in online environments, achieving a balance between GPU resource management and domain-specific classification accuracy. By prioritizing computationally intensive MLLM processes only for high-priority cases, this approach enhances the viability of MLLMs for real-time, high-stakes applications, such as video quality understanding tasks on TikTok.
By deploying COEF-VQ framework on TikTok internal video management plaform (VMP), our empirical study on two in-house tasks demonstrates the cost-efficiency of incorporating MLLM into a cascade framework and the superiority of our new cascade method on domain-specific classification tasks.
@@ 2.1. Multimodal LLM Architecture
The Multimodal LLM architecture, as illustrated in Figure 1, inherits the design of vision-language (VL) foundation model LLaVA-OneVision(Li et al., 2024) combined with audio encoder Whisper-Small(Radford et al., 2023). Several key components are integrated to process multimodal inputs, including a vision encoder for video frames, a VL modality alignment projector, a language model for both vision and text tokens, and an audio encoder for the video’s audio stream.
@@ 2.2.1. First Stage: Lightweight Model for Initial Filtering
In the first stage, a smaller, resource-efficient model is used to make initial predictions as soon as a video is posted. This lightweight model, often designed with fewer parameters and optimized for speed, performs a quick assessment of the content. Here, it assigns a preliminary classification score based on the likelihood of the video containing domain-specific content (such as potentially harmful content). Additionally, metadata such as video views (VV) or user engagement metrics can be incorporated to further refine the filtering criteria. This stage serves as a gatekeeper, quickly filtering out a significant number of videos with minimal resource consumption, and reducing the workload on the subsequent, more intensive stage.
@@ 2.2.3. Benefits and Implementation Considerations
This cascade deployment structure significantly lowers the GPU resource demands for online applications, making it possible to maintain scalability without incurring prohibitive costs. By using a smaller model for rapid pre-screening, the system reduces the number of inputs needing full MLLM processing, allowing for both cost efficiency and faster response times in high-traffic environments. Furthermore, this method ensures a balance between speed and accuracy, as the MLLM only engages when necessary. In terms of implementation, this cascade structure can be dynamically tuned; thresholds in the first stage (such as classification scores or engagement metrics) can be adjusted recurrently based on model performance, further optimizing resource allocation.
To leverage this newly proposed framework to keep users from inappropriate videos, we deploy COEF-VQ onto the Video Management Platform (VMP) at TikTok, such that newly published videos or updated contents could be interacted and safe-guarded through our cascade framework in real-time. Detailed introduction and explanation of VMP will be in the next section.
@@ 2.3. Deploy COEF-VQ onto Video Management Platform (VMP)
As a core component of Video Recommendation System, VMP plays a significant role in processing and gatekeeping videos published by creators, to ensure that the most engaging and appropriate content is presented to users. The functions and key components of VMP are illustrated in Figure 3. When a new video is published, model scores related to various quality and policy compliance factors, such as detecting inappropriate content on platforms like TikTok, are stored in the VMP stage. These scores are subsequently processed and utilized in downstream multi-stage recommendation funnel (like Recall, Pre-rank, Rank, etc.).
@@ 3. Experiments and Results
Table 1. Performance (in %) from baselines and MLLM on ICD and UCC experiments.
Task Approach Parameter Size F1 R@P60 R@P70 R@P80 R@P90
ICD    Base model 25M 59.1 57.6 44.1 27.9 20.2
   SIGLIP(Zhai et al., 2023) 679M 64.5 68.7 58.3 42.9 29.3
   MLLM (w/o audio) 8.3B 67.8 72.4 62.2 53.1 39.4
UCC    Base model 564M 77.6 89.8 84.5 74.8 54.1
   SIGLIP(Zhai et al., 2023) 699M 75.8 87.4 80.1 71.1 51.2
   MLLM 8.4B 80.5 92.3 87.8 79.3 62.2
Table 2. Performance (in %) of COEF-VQ on ICD and UCC experiments. QPS Ratio: query per second ratio required for calling MLLM service when video is published in TikTok
Task $P_{base}$ QPS Ratio R@P60 R@P70 R@P80 R@P90
ICD - 100 72.4 62.2 53.1 39.4
10 4.45 71.0 62.4 52.0 39.2
20 2.06 70.4 62.7 53.1 38.8
30 1.27 59.6 57.6 55.7 38.6
UCC - 100 92.3 87.8 79.3 62.2
50 16.7 91.2 86.9 79.8 62.6
60 13.3 89.6 85.4 79.9 62.3
70 10.8 88.4 83.1 77.7 62.8
@@ 3.1.1. Datasets
We pick two TikTok in-house video quality understanding tasks to test the effectiveness of our proposed MLLM and cascade structure. The two tasks are Inappropriate Content Detection(ICD) and Unoriginal Content Classification(UCC), respectively.
For ICD, it is a dataset aiming to determine whether the video violates the TikTok platform’s TnS policy or not, with 1 million training and validation videos and 100K test videos. For UCC, it aims to classify whether a posted video is UC(Unoriginal Content) or OC(Original Content). We use a training and validation dataset of 400K videos sampled from TikTok and 18K videos as the test set.
We SFT our MLLM based on each task’s training set and tune the hyper-parameters based on each task’s validation set. Both video quality tasks are evaluated in terms of classification F1 scores and recall under specific precisions.
@@ 3.1.2. Model Details
For MLLM, we invoke the LLaVA-OneVision model (Li et al., 2024) as the VL backbone, and consider Whisper-Small(Radford et al., 2023) as the audio encoder. We use ZERO2(Rajbhandari et al., 2020) and LoRA(Hu et al., 2021) to train our model. The rank and alpha used in LoRA is 32 and 64. For max frame length used in training, we respectively use 8 and 16 frames in ICD and UCC. We also rely on left padding when dealing with tokenization. For cascade structure, we directly use our online baseline models, which are ResNet(He et al., 2016) for ICD task, and a multi-tower multimodal model for UCC task including Swin-transformer(Liu et al., 2021), XLM-R(Conneau, 2019) and Whisper-Base(Radford et al., 2023).
@@ 3.2. Main Results
Comparisons between baselines and MLLM. In Table 1, we present a detailed comparison between MLLM and our baseline models on two classification tasks, demonstrating MLLM’s effectiveness. Notably, the baseline model employed for the Inappropriate Content Detection (ICD) task is a ResNet50, which operates on a single modality (visual features only); it’s a relatively low baseline and so we disable audio encoder in MLLM for ICD task for fairer comparison. Besides our own online baseline models described in the last section, we also include SIGLIP_Large(Zhai et al., 2023) for both tasks as another baseline of industry-as-usual. However, due to SIGLIP not including audio modality, its performance on UCC task is even below our online baseline for UCC.
@@
Results from COEF-VQ, the cascade structure. Table 2 provides a detailed analysis of our cascade structure’s performance across two classification tasks, highlighting its efficiency and precision. The cascade framework is designed to selectively trigger the MLLM stage based on predefined precision thresholds from the base model. This selective engagement optimizes resource usage by activating MLLM only when specific confidence criteria are met, allowing the base model to handle simpler cases and reserving MLLM’s capabilities for more ambiguous ones.
For tasks like Inappropriate Content Detection (ICD), where high recall is essential but often leads to lower precision (due to a greater risk of false positives), the cascade framework plays a critical role. By focusing MLLM’s processing power on reducing these false positives, the cascade system maximizes detection accuracy while controlling computational costs. Specifically, for the ICD task, the cascade system achieves performance nearly on par with MLLM alone but requires only 4.45% of the total query per second (QPS). This substantial efficiency gain is achieved without sacrificing performance, demonstrating the cascade’s cost-effectiveness at selectively deploying MLLM only when needed. Furthermore, for UCC task, we observe that under a threshold of Precision 60 for the base model, the cascade structure actually improves overall performance, achieving results even slightly superior to using MLLM exclusively with a Precision 80 threshold. This finding is notable because it indicates that our cascade approach not only conserves resources but also enhances performance by filtering out false positives with only 13.3% of the QPS directed towards MLLM. The cascade framework significantly reduces computational demands while maintaining accuracy in challenging high-recall, low-precision scenarios. By optimizing the interplay between the base model and MLLM, this approach provides an efficient and scalable solution that adapts well to variable workload demands.

### v2

@@ Abstract.
Recently, with the emergence of recent Multimodal Large Language Model (MLLM) technology, it has become possible to exploit its video understanding capability on different classification tasks. In practice, we face the difficulty of huge requirements for GPU resource if we need to deploy MLLMs online. In this paper, we propose COEF-VQ, a novel cascaded MLLM framework designed to enhance video quality understanding on the short-video platform while optimizing computational efficiency. Our approach integrates an entropy-based pre-filtering stage, where a lightweight model assesses uncertainty and selectively filters cases before passing them to the more computationally intensive MLLM for final evaluation. By prioritizing high-uncertainty samples for deeper analysis, our framework significantly reduces GPU usage while maintaining the strong classification performance of a full MLLM deployment. To demonstrate the effectiveness of COEF-VQ, we deploy this new framework onto the video management platform (VMP) at the short-video platform, and perform a series of detailed experiments on two in-house tasks related to video quality understanding. We show that COEF-VQ leads to substantial performance gains from the offline evaluation in these two tasks and effectively enhances platform safety with limit resource consumption, significantly reducing inappropriate content video view rate by 9.9% in a online A/B test without affecting engagement. Post-launch monitoring confirmed sustained improvements, validating its real-world impact.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/3711896.3737203ccs: Information systems Recommender systemsccs: Computing methodologies Artificial intelligenceccs: Information systems Language models
Figure 1. Illustration of our proposed unified Multimodal LLM architecture: video frames and text tokens are processed and early-fused through a VL foundational model, and audio signal is processed through an Audio encoder and then late-fused with the hidden representations from VL model.
@@ 1. Introduction
Therefore, in this paper we propose COEF-VQ, which is composed of two key novelties:
@@
1)1)1 ) we first design a unified Multimodal LLM framework integrating video frames, texts, and audio information, which goes beyond the existing multi-tower baseline deployed on our platform (Conneau, 2019; He et al., 2016; Liu et al., 2021; Radford et al., 2023) and typical industry practices (Zhai et al., 2023). This enhanced model architecture allows for more sophisticated feature fusion and deeper cross-modal interactions, boosting the video classification performance. By doing so, it alleviates the limitations of traditional multi-tower structures, making the model more sensitive and accurate in discerning domain-specific tasks.
@@
However, despite the powerful capabilities of MLLM, the large number of parameters in MLLM, which enables them to perform robust multimodal reasoning, requires substantial GPU memory and processing speed. This demand, when scaled for online deployment on short-video platforms, would lead to considerable GPU resource consumption, affecting both deployment costs and latency in real-time, online video service.
2)2)2 ) we develop a cascaded serving framework for online deployment, designed to optimize resource utilization while maintaining strong performance. In the first stage, a lightweight model (typically the baseline) generates initial predictions at the time of video posting. To enhance filtering efficiency, we introduce an entropy-based criterion, where videos with low uncertainty are filtered out, significantly reducing the volume passed to the next stage. This ensures that only uncertain cases undergo further evaluation. In the second stage, an MLLM processes the remaining videos with higher accuracy, effectively reducing false positives. This entropy-driven pre-filtering not only minimizes unnecessary computational overhead but also ensures that the MLLM focuses on the most ambiguous or critical cases, improving classification reliability. By prioritizing MLLM processing for high-uncertainty inputs, our approach achieves a scalable and cost-effective deployment of MLLMs in real-time environments. This framework is particularly well-suited for high-stakes applications, such as video quality understanding on short-video platforms, where balancing resource efficiency with classification accuracy is crucial.
By deploying COEF-VQ framework on internal video management plaform (VMP) from the short-video platform, our empirical study on two in-house tasks demonstrates the cost-efficiency of incorporating MLLM into a cascade framework and the superiority of our new cascade method on domain-specific classification tasks. We also conduct a series of experiments on Inappropriate Content Detection (ICD) for online deployment, a key task in maintaining platform safety. Our model deployment reduces GPU usage to only 5% of a full-traffic MLLM deployment during online A/B testing. Over a 10-day online experiment with 10% traffic allocation per group, our framework significantly reduced the inappropriate content video view rate (ic_vvr) by 9.9% without impacting core engagement metrics. The post-launch evaluation, spanning more than a month of global monitoring, further confirmed the effectiveness of COEF-VQ, demonstrating sustained reductions in ic_vvr and validating its real-world impact.
@@ 2.1. Multimodal LLM Architecture
The Multimodal LLM architecture, as illustrated in Figure 1, inherits the design of vision-language (VL) foundation model combined with audio encoder. Several key components are integrated to process multimodal inputs, including a vision encoder for video frames, a VL modality alignment projector, a language model for both vision and text tokens, and an audio encoder for the video’s audio stream.
@@ 2.2.1. First Stage: Lightweight Model for Initial Filtering
In the first stage, a smaller, resource-efficient model is used to make initial predictions as soon as a video is posted. This lightweight model, often designed with fewer parameters and optimized for speed, performs a quick assessment of the content. Here, it assigns a preliminary classification score based on the likelihood of the video containing domain-specific content (such as potentially harmful content). To prioritize informative and ambiguous samples, we propose an entropy-based filtering criterion in our cascade structure. We compute the entropy score H, defined as:
(5) $\textbf{H}(p)=-\sum_{i=1}^{N}p_{i}\log p_{i}$
where $p_{i}$ represents the predicted probability for class $i$ in an $N$-class classification task. Videos with low entropy (i.e., confident predictions) are filtered out efficiently, while those with high entropy indicating uncertain or ambiguous cases, are sent to the next stage for further analysis. Additionally, metadata such as video views (VV) or user engagement metrics can be incorporated to further refine the filtering criteria. This stage serves as a gatekeeper, quickly filtering out a significant number of videos with minimal resource consumption, and reducing the workload on the subsequent, more intensive stage.
@@ 2.2.3. Benefits and Implementation Considerations
This cascade deployment structure significantly lowers the GPU resource demands for online applications, making it possible to maintain scalability without incurring prohibitive costs. By using entropy for effective pre-screening, the system reduces the number of inputs needing full MLLM processing, allowing for both cost efficiency and faster response times in high-traffic environments. Furthermore, this method ensures a balance between speed and accuracy, as the MLLM only engages when necessary. In terms of implementation, this cascade structure can be dynamically tuned; thresholds of entropy score in the first stage can be adjusted recurrently based on model performance, further optimizing resource allocation.
@@ 2.3. Deploy COEF-VQ onto Video Management Platform (VMP)
To leverage this newly proposed framework to keep users from inappropriate videos, we deploy COEF-VQ onto the Video Management Platform (VMP), such that newly published videos or updated contents could be interacted and safe-guarded through our cascade framework in real-time.
As a core component of Video Recommendation System, VMP plays a significant role in processing and gatekeeping videos published by creators, to ensure that the most engaging and appropriate content is presented to users. The functions and key components of VMP are illustrated in Figure 3. When a new video is published, model scores related to various quality and policy compliance factors, such as detecting inappropriate content on platforms, are stored in the VMP stage. These scores are subsequently processed and utilized in downstream multi-stage recommendation funnel (like Recall, Pre-rank, Rank, etc.).
@@ 3. Experiments and Results
Table 1. Performance (in %) from baselines and MLLM on ICD and UCC experiments. MMM: Multi-tower Multimodal Model. The maximum BETA variance across all models for ICD is under 0.08 (in %), while for UCC is under 0.01 (in %).
Task Approach Parameter Size F1 R@P50 R@P60 R@P70 R@P80 R@P90
ICD    ResNet50(He et al., 2016) 25M 59.1 65.1 57.6 44.1 27.9 20.2
   SIGLIP(Zhai et al., 2023) 679M 64.5 75.6 68.7 58.3 42.9 29.3
   Qwen2.5_VL(Bai et al., 2025) 8.3B 66.5 80.3 72.3 60.2 53.7 39.8
   MLLM (w/o audio) 8.3B 67.8 81.1 72.4 62.2 53.1 39.4
UCC    MMM 564M 77.6 93.2 89.8 84.5 74.8 54.1
   SIGLIP(Zhai et al., 2023) 699M 75.8 91.3 87.4 80.1 71.1 51.2
   Qwen2.5_VL(Bai et al., 2025) 8.4B 79.5 93.4 91.9 86.9 78.9 61.7
   MLLM 8.4B 80.5 95.8 92.3 87.8 79.3 62.2
Table 2. Performance (in %) of COEF-VQ on ICD and UCC experiments. QPS Ratio: query per second ratio required for calling MLLM service when video is published in the short-video platform. H_base: entropy score threshold from the base model in first stage used for filtering.
Task $H_{base}$ QPS Ratio R@P50 R@P60 R@P70 R@P80
ICD - 100 81.1 72.4 62.2 53.1
0.3 5.19 81.5 73.4 62.6 52.0
0.6 2.22 81.7 75.0 65.2 43.9
0.9 0.79 79.1 71.9 58.3 43.1
UCC - 100 95.8 92.3 87.8 79.3
0.3 16.7 94.6 92.2 87.3 79.3
0.6 10.0 93.5 91.1 86.7 79.3
0.9 4.55 93.0 89.9 85.8 76.5
@@ 3.1.1. Datasets
We pick two in-house video quality understanding tasks to test the effectiveness of our proposed MLLM and cascade structure. The two tasks are Inappropriate Content Detection(ICD) and Unoriginal Content Classification(UCC), respectively. For ICD, it is a dataset aiming to determine whether the video violates the short-video platform’s TnS policy or not, with 1 million training and validation videos and 100K test videos. For UCC, it aims to classify whether a posted video is UC(Unoriginal Content) or OC(Original Content). We use a training and validation dataset of 400K videos sampled from the short-video platform and 18K videos as the test set. We SFT our MLLM based on each task’s training set and tune the hyper-parameters based on each task’s validation set.
@@ 3.1.2. Model Details
For MLLM, we invoke the LLaVA-OneVision with 7B model (Li et al., 2024) as the VL backbone, and consider Whisper-Small(Radford et al., 2023) as the audio encoder. We use ZERO2(Rajbhandari et al., 2020) and LoRA(Hu et al., 2021) to train our model. The rank and alpha used in LoRA is 32 and 64. For max frame length used in training, we respectively use 8 and 16 frames in ICD and UCC. We also rely on left padding when dealing with tokenization. For cascade structure, we use the best baseline models as the base model, which are SIGLIP_Large(Zhai et al., 2023) for ICD task, and a multi-tower multimodal model for UCC task including Swin-transformer(Liu et al., 2021) pretrained by in-house data, XLM-R(Conneau, 2019) and Whisper-Base(Radford et al., 2023).
3.1.3. Offline Evaluation Metrics
Both video quality tasks are evaluated in terms of classification F1 scores, recall at various precision and beta variance.
Recall at various Precision corresponds to the recall rate for binary classification of whether the label is ”inappropriate” for ICD task, or ”unoriginal” for UCC task.
Beta Variance is defined as the largest variance of a Beta distribution used to model the posterior uncertainty of the positive-class proportion (Sensoy et al., 2018). Concretely, if a subset of samples has $\alpha$ for true positives amount + 1 and $\beta$ for false positives amount + 1, we define variance of a Beta distribution as
(6) $\text{Variance}=\frac{\alpha\beta}{(\alpha+\beta)^{2}(\alpha+\beta+1)}$
Then compute this variance for each precision/recall subset and take the maximum as the maximum Beta variance, reflecting the highest level of uncertainty in estimating the true positive rate. We consider that maximum Beta Variance below 1% suggests strong certainty.
@@ 3.2. Main Results
Comparisons between baselines and MLLM. In Table 1, we present a detailed comparison between MLLM and our baseline models on two classification tasks, demonstrating MLLM’s effectiveness. Notably, ResNet50(He et al., 2016) used for Inappropriate Content Detection (ICD) task operates on a single modality (visual features only); it’s a relatively low baseline and so we disable audio encoder in MLLM for ICD task for fairer comparison. Besides ResNet50, we also include SIGLIP_Large(Zhai et al., 2023) and Qwen2.5_VL(Bai et al., 2025) for both tasks as baselines of industry-as-usual. However, due to SIGLIP not including pretrained in-house video knowledge, its performance on UCC task is even below our internal multi-tower multimodal model for UCC.
@@
Results from COEF-VQ, the cascade structure. Table 2 provides a detailed analysis of our cascade structure’s performance across two classification tasks, highlighting its efficiency and precision. The cascade framework is designed to selectively trigger the MLLM stage based on predefined entropy thresholds from the base model. This selective engagement optimizes resource usage by activating MLLM only when specific entropy criteria are met, allowing the base model to handle simpler cases and reserving MLLM’s capabilities for more ambiguous ones.
For tasks like Inappropriate Content Detection (ICD), where high recall is essential but often leads to lower precision (due to a greater risk of false positives), the cascade framework plays a critical role. By focusing MLLM’s processing power on reducing these false positives, the cascade system maximizes detection accuracy while controlling computational costs. Specifically, for ICD task, we observe that above an entropy threshold of 0.6 for the base model, the cascade structure actually improves overall performance, achieving results even slightly superior to using MLLM exclusively with a Precision 70 threshold. This finding is notable because it indicates that our cascade approach not only conserves resources but also enhances performance by filtering out false positives with only 2.22% of the QPS directed towards MLLM. Furthermore, for the UCC task, the cascade system achieves performance nearly on par with MLLM alone with 16.7% of the total query per second (QPS). This substantial efficiency gain is achieved without sacrificing performance, demonstrating the cascade’s cost-effectiveness at selectively deploying MLLM only when needed.
The cascade framework significantly reduces computational demands while maintaining accuracy in challenging high-recall, low-precision scenarios. By optimizing the interplay between the base model and MLLM, this approach provides an efficient and scalable solution that adapts well to variable workload demands.
@@
Table 6. Performance (in %) between Confidence-based filtering approach and Entropy-based filtering approach with 10% QPS Ratio.
Approach R@P50 R@P60 R@P70 R@P80
     Confidence-based 92.9 88.4 83.1 77.7
     Entropy-based 93.5 91.1 86.7 79.3